### v2

# How individual vs shared coordination governs the degree of correlation in rotational vs residence times in a high-viscosity lithium electrolyte

Abstract
Commercially used carbonate-based electrolytes in lithium-ion batteries are susceptible to many challenges, including flammability, volatility, and lower thermal stability. Solvated ionic liquids of LiTFSI salt (lithium bis(trifluoromethylsulfonyl)-amide) and glyme-based solvents are potential alternative candidates for commonly used electrolytes. We perform classical molecular dynamics (MD) simulations study the effect of concentration and temperature on the translational and rotational dynamics. The radial distribution function shows stronger coordination of Li+ ions with tetraglyme(G4), as shown in earlier studies, and forms a stable [Li(G4)]+ cation complex. The self-diffusion coefficients are lower than the values experimentally observed but show better improvement over other classical force fields. An increase in the salt concentrations leads to a higher viscosity of the system and reduces the overall ionic mobility of Li+ ions. Diluting the system with a larger number of glyme molecules leads to shorter rotational relaxation times for both TFSI and tetraglyme. Ion-residence times show that Li+ ions form stable and long-lasting complexes with G4 molecules than TFSI anions. The residence time of [Li(G4)]+ complex increases at higher salt concentrations due to the availability of fewer G4 molecules to coordinate with a Li+ ion. G4 is also seen to form polydentate complexes with Li+ without a shared coordination, allowing rotation without breaking coordination, unlike TFSI, which requires coordination disruption for rotation. This distinction explains the poor correlation between rotation and residence time for G4 and the strong correlation for TFSI.
I Introduction
The growing demand for clean energy across the world requires efficient energy storage devices. Batteries and supercapacitors are among the widely used electrochemical devices for energy storage. Lithium-ion batteries (LIBs) have made significant progress in the past three decades in reducing their overall cost with an increase in energy density. The cost has decreased from $1,000 per kWh in 2009 to $180 per kWh in 2020, while the energy density has increased from 125 Wh/kg to 250 Wh/kg. The LIB market was $36.7 billion in 2019 and is expected to reach $128.3 billion by 2027, with an estimated compound annual growth rate of 18% from 2020 to 2027.Feng et al. (2022); Nitta et al. (2015). The electrolyte is an important component of batteries that determines the cyclic efficiency and the charge-discharge rate of a battery. Generally, ideal electrolytes should possess high ionic conductivity and good chemical and thermal stability. Commercialized LIBs use electrolytes of fluoride salts and organic solvents, which provide high ionic conductivity but are susceptible to leakages and flammability and are responsible for most fire accidents in electric vehicles. Various alternative electrolyte systems, such as ionic liquids (ILs),Rana, Thakur, and Dosanjh (2023) solvated ionic liquids (SILs),Nachaki and Kuroda (2024) inorganic solid electrolytes,Yu et al. (2025) soft-solid electrolytes,Prakash et al. (2023) and composite electrolytes,Zhang et al. (2023) have been proposed to increase the safety aspects and energy density of the batteries. SILs are considered an alternative to conventional ILs owing to their unique structural characteristics and lower viscosity. SILs are formed when a cation (e.g., Li+ here) forms a coordinated complex with the donor sites of the solvent molecule, with anions being solvent-separated and occupying the interstitial volume. The glyme family of solvents, specifically those with higher chain lengths, are a better alternative to organic solvents in SILs due to their lower flammability, less toxicity, and higher thermal stability. In SILs, the currently used higher chain length solvents are triglyme (G3), tetraglyme (G4), and pentaglyme (G5) with LiTFSI and LiPF6 salts. LiTFSI salt shows excellent thermal and chemical stability up to 360 °C and dissociates easily in various types of solvents.Wang et al. (2019); Di Lecce et al. (2022); Lingua et al. (2024); Rajendran, Ganesan, and Arava (2024); Dong and Bedrov (2018)
Higher chain-length glyme molecules have better physiochemical properties compared to shorter chain-length glyme molecules. Equimolar concentrations of triglyme (G3) and tetraglyme (G4) have been investigated with lithium salts using various experimental (Raman spectroscopy, Small angle X-Ray Scattering, Diffusion nuclear magnetic resonance) and computational techniques.Aguilera et al. (2015); Yoshida et al. (2011); Tamura et al. (2010) Structural properties like radial distribution function (RDF) and radius of gyration () are useful in predicting the local solvation shell and compactness of atoms or molecules, respectively. The local structure of Li+ ions prefers to remain in a five-fold (or four-fold) coordination with oxygen of G4 and TFSI combined, where the major contribution comes from G4.Tsuzuki et al. (2014); Callsen et al. (2017); Sun and Hamada (2018) Based on G4 or TFSI coordination, the LiTFSI/G4 complexes are classified as contact ion pairs (CIP) and solvent-separated ion pairs (SSIP). In CIP, oxygen atoms on TFSI directly coordinate with solvated Li+ ions, whereas in SSIP, TFSI anions interact with the [Li(G4)]+ complex. The SSIP configurations were found to be energetically more favorable compared to CIP, stating that Li+ ions prefer to remain as a [Li(G4)]+ complex in the system.Aguilera et al. (2015) Experimental techniques like 2D IR spectroscopy, along with molecular mechanics-based non-reactive classical molecular dynamics (MD) simulations and first principle calculations, have been used to assess and characterize the distribution of these various structural complexes.Rajput et al. (2015); Callsen et al. (2017); Rushing, Leonik, and Kuroda (2019); Galvez-Aranda and Seminario (2021); Chen et al. (2024); Tan et al. (2024) However, the challenge with classical MD is that it does not capture the structural changes in solvated structures during the dynamics very well compared to Ab initio (AIMD) simulations. Still, inaccessibility to longer timescales in AIMD simulations forces the use of classical models, especially for longer time-correlated properties like translational and rotational diffusion. Experimentally, the diffusion coefficients for Li+ , TFSI, and G4 are calculated at 30 °C and 80 °C (discussed later in the subsection Translational dynamics).Tamura et al. (2010); Yoshida et al. (2011); Zhang et al. (2014); Harte et al. (2023) Using MD simulations with a polarizable force field, Dong et al.Dong and Bedrov (2018) found that the diffusion coefficients of species at 30 °C and 80 °C are one or two order of magnitude less than the experimental diffusion coefficients. Other research groups also used non-polarizable OPLS-force fields with scaled charges (to include polarization effects) and found similar diffusion coefficients at 227 °C.Thum et al. (2020) No experimental study has been done so far that highlights the system dynamics at such elevated temperaturesHarte et al. (2023). The use of scaled charges can model the structure right by over-simplification of the many-body interactions. Thus, the use of more robust methods to compute the partial charge of molecules is required. Dong’s group also calculated the lifetime of the Li+ ions with TFSI and G4 at 373 K, which depends upon the salt concentration. A smaller concentration of salt leads to a greater number of free glyme molecules in the system, which goes through a faster exchange event with glyme molecules binding with Li+ ions, thus reducing the lifetime of the .Dong and Bedrov (2018) The translation diffusion, exchange events, and residence times also dictate the coiling and bending of geometries of G4 and TFSI, respectively. However, no modeling of rotational dynamics has been conducted yet that shows the effect of the coiling/bending of one molecule on another or on the rotational lifetime of molecules. The rotational contribution is an important factor in deciding the lifetime of big continuous clusters of Li+ ions with TFSI and G4.
This study investigates the extreme range of LiTFSI and G4 mixture compositions (from 1:4 to 2:1) at various temperatures with classical MD simulations using a modified OPLS force-field.Jorgensen, Maxwell, and Tirado-Rives (1996) The partial charges for atoms of LiTFSI and G4 are reparametrized here (Fig. 1). A correct structure prediction and long timescale simulations performed in the present work provide a reliable set of results for diffusion coefficients, residence times, rotational relaxation times, and cluster sizes and theorize a not-yet-explored interplay of these.
II Computational Details
Four mixtures with different LiTFSI/G4 ratios (1:4, 1:2, 1:1, and 1:4) were studied, as shown in Table S1 of supplementary information. These mixtures were modeled using a charge-reparameterized OPLS force field. Partial atomic charges for the LiTFSI and G4 molecules were calculated (Fig. 1a) using a range-separated hybrid functional wB97x-D4 functionalChai and Head-Gordon (2008); Caldeweyher, Bannwarth, and Grimme (2017) with a sufficiently large basis set def2-tzvpp, utilizing the ORCA 5.0.1 packageNeese (2022). The PMEEssmann et al. (1995) algorithm was used to compute the long-range electrostatic interactions. The cut-off for Lennard-Jones and electrostatic interactions is set to 10 Å. A time-step of 1 femtosecond is used throughout the simulations. All simulations were conducted using the Gromacs 2021 Abraham et al. (2015) software package. For post-processing and analysis, VMD 1.9.3 Humphrey, Dalke, and Schulten (1996), TravisBrehm et al. (2020), and in-house scripts developed in python were used.
The different ratio mixtures were initially minimized using the conjugate-gradient algorithm. Each system was heated from 300 K to 600 K with a heating rate of 10 K/ns to mix it well. A modified set of Berendsen thermostatBussi, Donadio, and Parrinello (2007) and barostatBerendsen et al. (1984) were used to manage fluctuations in temperature and pressure, respectively. As the systems were heated to a particular temperature at successive intervals, an equilibration run was performed in NVT ensemble conditions for 10 ns. Fig. 1(b-e) shows post-equilibrated mixtures at 300 K. Following this, the systems were cooled again from 600 K to 300 K at the interval of 50K/ns with 10 ns of equilibration at each particular temperature. Final production runs were performed at 300 K and 500 K in NVT ensemble conditions for 100 ns to obtain sufficiently long trajectories for analysis.
III Results and discussion
III.1 Solvation structure
The interatomic interactions of TFSI and G4 around Li+ ions are calculated as radial distribution functions (RDF):
where is RDF, is the number of atoms in the shell, is the mean atom-density, and is the width of the spherical shell inside which the number of atoms is considered. The coordination number is calculated by integrating the RDF over the chosen spherical shell around the central atom. The neutron scattering confirms that in an equimolar mixture of LiTFSI and G4, 95 % G4 molecules are coordinated by Li+ ions and the remaining 5 % have an incomplete Li…O(G4) coordination.Murphy et al. (2016) The experimental peaks are observed at 1.9 Å for Li…O(G4), and 2.0 Å for Li…O(TFSI) in an equimolar (1:1) mixture.Murphy et al. (2016) Our simulations are in a good agreement with these observations as the first solvation shell of a Li+ ion is seen to form at 1.9 Å with O atoms of G4 and at 2.0 Å with O atoms of TFSI (Fig. 2). From the integration of these RDFs (within a cut-off radius of 2.3 Å), the coordination number in the equimolar ratio at 300 K is 3.6 for Li…O(G4) and 0.7 for Li…O(TFSI). The variation in the coordination number of Li+ ions is observed at different salt concentrations. With the increase in the concentration of glyme molecules, the coordination number of Li…O(G4) increases nearly to 4 at 1:2 and 1:4 ratios. Also, the increase in the concentration of solvent leads to an increase in the intensity of RDF of the Li…O(G4), whereas the intensity of RDF of Li…O(TFSI) decreases as more number of Li+ ions find free G4 and get attached to them. Changing the concentration of salt and solvent does not affect the radius of the first coordination shell around lithium.
The coordination number of Li…O(G4) in equimolar ratio decreased to 3.3 (500 K) from 3.6 (300 K). This could happen due to the accelerated dynamics, which leads to a faster exchange between G4 molecules at higher temperatures within the solvation shell of lithium. This also leads to a decrease in the lifetime of Li…O(G4) complex which is confirmed through ion-pair/ion-solvent existence autocorrelation function (discussed vide infra in subsection E). Also, the intensity of RDF peaks showed a small decrease at 500 K for both Li…O(TFSI) and Li…O(G4) when compared to RDF at 300 K.
III.2 Radius of gyration
| Temp (300 K) | Temp (500 K) | ||||
| Species | Average | Std. dev. | Average | Std. dev. | |
| 1:4 | TFSI | 0.2564 | 0.0006 | 0.2568 | 0.0006 |
| G4 | 0.3808 | 0.0016 | 0.3886 | 0.0018 | |
| 2:1 | TFSI | 0.2540 | 0.0002 | 0.2550 | 0.0003 |
| G4 | 0.3628 | 0.0007 | 0.3672 | 0.0034 | |
Since both the TFSI anion and G4 are long molecular structures, their coiling leads to a significant variation in the radii of gyration (). The of these species is an important factor that is sensitive to the temperature and concentration of electrolyte mixture. We used MD trajectories to calculate for TFSI anions and G4 molecules for each ratio and at 300 K and 500 K (Fig. S1 and S2 in supplementary information). As shown in Table 1, the of TFSI and G4 in the most diluted system (1:4) at 300 K is, 0.256 nm and 0.381 nm, respectively. At a higher concentration, 2:1, the for both TFSI and G4 decreases to 0.254 nm and 0.363 nm, suggesting a more coiled configuration, particularly for G4, which shrinks by 5 %. The for both TFSI and G4 decrease with concentration of LiTFSI. This indicates that at higher salt concentrations, most G4 molecules coordinate individually with the Li+ ions, thus increasing their overall compactness. At lower salt concentrations, there are more free G4 molecules available to share the coordination leading to higher . As expected at higher temperature, for TFSI and G4 increase from 0.2564 nm to 0.2568 nm and from 0.3808 nm to 0.3886 nm respectively, when increasing the temperature from 300 K to 500 K. Even these slight changes in is an important factor which can lead to an interplay in translational and rotational dynamics and ion-residence times in these mixtures (discussed vide infra).
III.3 Translational dynamics
| Temp (K) | |||
| aExperiments - Tamura et al.Tamura et al. (2010) | |||
| 303 | 1.26 | 1.22 | 1.26 |
| aExperiments - Yoshida et al.Ueno et al. (2012) | |||
| 303 | 1.31 | 1.22 | 1.29 |
| aExperiments - Zhang et al.Zhang et al. (2014) | |||
| 303 | 1.26 | 1.22 | 1.26 |
| aExperiments - Harte et al.Harte et al. (2023) | |||
| 303 | 1.01 | 1.04 | 0.87 |
| 353 | 27.90 | 7.08 | 6.33 |
| bMD Simulations - Dong et al.Dong and Bedrov (2018) | |||
| 303 | 0.41 | 0.40 | 0.41 |
| cMD Simulations - Tsuzuki et al.Tsuzuki et al. (2014) | |||
| 403 | 0.78 | 0.73 | 0.78 |
| cMD Simulations - Shinoda et al.Shinoda et al. (2018) | |||
| 503 | 4.5 | 5.7 | 4.6 |
| dMD Simulations - Thum et al.Thum et al. (2020) | |||
| 500 | 53 | 59 | 54 |
| eMD Simulations - Harte et al.Harte et al. (2023) | |||
| 500 | 52.02 | 56.63 | 53.10 |
| 700 | 96.16 | 113.12 | 96.69 |
| fMD Simulations - This work | |||
| 300 | 0.02 | 0.04 | 0.03 |
| 500 | 12.48 | 23.70 | 12.38 |
| aPFG-NMR | |||
| bMD simulation, polarizable APPLE&P force field | |||
| cMD simulation, OPLS-AA and CL&P force field, no charge scaling | |||
| dMD simulation, OPLS-AA and CL&P force field, with scaled charges | |||
| eMD simulation, OPLS-AA force field, with scaled charges | |||
| fMD simulation, OPLS-AA force field, no charge scaling | |||
The translational diffusion of ions and G4 is ascertained with the calculation of mean square displacement (MSD) for all trajectories. Since Li+ ions are strongly coordinated with O atoms of TFSI and G4, at high-concentration (e.g., 2:1) and low-temperature (e.g. 300 K), their movement (Li+ ions) is heavily vibrational in their coordination ’cage’. The trajectory lines for 2:1 and 1:4 compositions (Fig. 3) at 300 K and 500 K show that for 2:1 at 300 K (Fig. 3a), all Li+ ions move in the coordination cage with a vibrational motion of 0.1 - 0.3 Å of amplitude between two frames separated over 5 ps time-interval. This effect is well-known in other inorganic and soft-solid electrolytes. The self-diffusion coefficients () are obtained from the diffusive regime of MSD vs time plots (refer to Fig. S3 and S4 in supplementary information), utilizing the Einstein relationship as:
Here, is the position of th atom at time , is the lag time, is the dimensionality of the system, is the number of particles, and is the mean square displacement average over multiple times origin. The calculated values of self-diffusion coefficients at 300 K and 500 K are reported in Table S2 of supplementary information, and provided in Fig. 4 (at 500 K). Experimentally, the self-diffusion coefficient of Li+ ions () is measured using both impedance spectroscopy and pulsed-field gradient NMR (Table III.3). Yoshida et al. (2011); Zhang et al. (2014); Harte et al. (2023) Although the use of scaled charges, which is a brute-force method to get the appropriate local interactions and diffusion of Li+ ions in these systems, predicts better agreement as shown in other works.Thum et al. (2020); Harte et al. (2023) The diffusion coefficients of G4 are similar to that of Li+ ions for a 1:1 system (12.38), showing the formation of a stable [Li(G4)]+ complexes. This suggests that lithium primarily moves in the form of the [Li(G4)]+ complex. With more G4 molecules available in the mixture, , and increase due to a faster exchange between G4 molecules inside Li+ ion solvation shell. However, increasing the concentration of LiTFSI leads to an increase in the pairing probability of [Li(G4)]+ and TFSI anion, causing sluggish diffusion of both the ions.
III.4 Rotational dynamics
The rotational autocorrelation function (RACF) is employed to analyze the orientational dynamics of TFSI and G4. In the case of TFSI anion, a specific atom triplet is selected, comprising S-N-S atoms. A vector is defined by taking the cross-product of the two vectors connecting the S-N and N-S atoms. Similarly, for the G4 molecule, the atom triplet consists of C-O-C atoms, where O is the central oxygen atom of the chain. The RACF for the vector is written as
where and are the vectors chosen for molecules at time t = and t = .
Fig. S5 in supplementary information shows that for the system with the highest concentration of G4 molecules (1:4), RACF for both TFSI and G4 decays slower with increasing salt concentration. In dilute mixtures, larger number of free G4 molecules in the system are partially/not coordinated with Li+ ions, resulting in fewer obstructions in their rotational dynamics. With increasing salt concentration, the rotational decay becomes much slower, due to the higher viscosity of the system and increase in the shared coordination of TFSI and G4 with Li+ ions. Also, G4 molecules rotate more slowly than TFSI anions, primarily due to a bigger radius of gyration and also due to their higher coordination with Li+ ions.
The rotational relaxation time (Table 3) is calculated as
At 300 K, the C(t) for some compositions did not decay sufficiently during the simulation time; hence for those is calculated analytically here by fitting C(t) with two exponential functions. At 300 K, is 295 ns for TFSI and 893 ns for G4 for 2:1 mixture. As the amount of G4 increases, decreases sharply for both TFSI and G4. The value of at 300 K is in the order of ns and decreases to the order of ps as the temperature is raised to 500 K. The local structure and the nature of coordination of Li+ ions alters for a couple of orders across these concentrations. An analysis on the correlation of with the residence time of G4…Li and TFSI…Li and formation of interatomic networks (clusters) is discussed further to understand this complexity.
| Temp (300 K) | Temp (500 K) | |||
| 1:4 | 0.197 | 6.017 | 0.005 | 0.0265 |
| 1:2 | 0.591 | 22.587 | 0.006 | 0.0588 |
| 1:1 | 15.698 | 189.146 | 0.022 | 0.164 |
| 2:1 | 295.22 | 893.312 | 0.264 | 0.589 |
III.5 Ion-residence times
The ion-residence times are calculated for Li+ ions with TFSI and G4 as the duration which a Li+ ion would spend with a TFSI anion or a G4 molecule before completely breaking away from the interaction. The residence time of Li+ aggregates formed with TFSI and G4 is computed as an ion-pair (or ion-solvent) existence autocorrelation function (IEAF), which can be written as an auxiliary function
and the IEAF for is
The residence time of the aggregates formed can be computed by twice the integral of IEAF
for Li…O(G4) and Li…O(TFSI) interactions is calculated within the first solvation shell of lithium. The radius of the solvation shell is taken from the first minima of the corresponding RDFs. Fig. S6 in supplementary information shows the decay of IEAF for Li-O, for both TFSI and G4 at 300 K and 500 K. The decay is too sluggish for Li…O(TFSI) and Li…O(G4) in the systems with higher salt concentration. For all these systems is calculated numerically (for well-decayed) and analytically (for sluggish-decayed) tabulated in Table S3 on the supplementary information. The value of is 6 s for Li…TFSI and 30 s for Li…G4 for 2:1 mixture at 300 K, which decreases to 1 ns and 882 ns, respectively, as the system is diluted with more G4 (1:4). Dong et al.Dong and Bedrov (2018) calculated the residence time of Li…O(TFSI) at various temperatures and showed that, for the 1:1 system, the residence time for Li…O(TFSI) is 7 ns at 303 K and decreases with increasing the concentration of free G4 molecules and at higher temperatures. In our study, for equimolar system, the residence time of Li…O(TFSI) is 445 ns at 300 K and decreased to 1 ns at 500 K, as shown in Table S3 in the supplementary information. Similarly, the residence time of Li…O(G4) is 34 s at 300 K and decreases by three orders to 46 ns at 500 K. In a similar study for an equimolar LITFSI/G4, Shinoda et al.Shinoda et al. (2018) reported that the residence time of Li…TFSI and Li…G4 at 503 K was 2.5 ns and 87 ns, respectively. More dilution leads to an increase in the number of free G4 molecules in the system, enabling their rapid exchange with the Li+ ion coordination shell, subsequently decreasing the residence time.
We propose that such long-residence times, particularly at high concentration and low temperature, could affect the rotational motion of these entities with a significantly large . If an ion-solvent or ion-pair network has a long lifetime (), then their would also be proportionately longer. To probe this, we computed the correlation of and for each G4 and TFSI, at 300 K and 500 K. Here each dataset has four pairs of values for the different concentrations of G4 and TFSI. Plotted on a log-log scale (Fig. 5), lifetimes (rotational-individual and residence with Li) of TFSI anion are fitted as and at 300 K and 500 K, respectively - exhibiting a greater than 99% correlation in both cases of temperature. However, for G4, there is surprisingly low-correlation of 66% (at 300 K) and 89% (at 500 K) between the lifetimes. It can be understood that an increase from 300 K to 500 K in temperature simply aligns the rotational motion with the residence time better, but there is a stark contrast of this correlation with TFSI. To investigate further, we analyzed ionic aggregates, most-probable clusters and their nature below.
III.6 Ionic aggregates and clusters
| 300K | 500K | |||||||
| Li-TFSI | Li-G4 | Li-TFSI | Li-G4 | |||||
| Uncoord. | Size of Max. | Uncoord. | Size of Max. | Uncoord. | Size of Max. | Uncoord. | Size of Max. | |
| 1:4 | 5.675 | 187.3013 | 0.139 | 539.8607 | 5.285 | 172.5438 | 0.405 | 539.5926 |
| 1:2 | 7.635 | 415.0555 | 1.708 | 645.6169 | 8.363 | 412.1730 | 2.282 | 644.5518 |
| 1:1 | 3.000 | 396.9527 | 0.375 | 455.5046 | 2.790 | 435.9582 | 0.850 | 327.9481 |
| 2:1 | 11.454 | 852.4622 | 4.004 | 631.0744 | 8.575 | 845.2045 | 3.002 | 637.8953 |
| Temp (300 K) | Temp (500 K) | |||
| Single | Maximum | Single | Maximum | |
| 1:4 | 77.6 | 569 | 77.806 | 553 |
| 1:2 | 147.762 | 712 | 122.905 | 705 |
| 1:1 | 91.532 | 637 | 65.195 | 634 |
| 2:1 | 5.857 | 1070 | 9.58 | 1062 |
All the concentrations of LITFSI/G4 are observed to form long networks where Li, TFSI and G4 coordinate in a shared and unshared manner to form various clusters. Fig. 6 shows the dynamic bonds around each Li+ ion in each case of concentration at 300 K. The snapshots show that for the dilute mixtures (1:4 and 1:2), most of these clusters are individual, and/or unshared (Fig. 6 a, b). For the concentrated mixtures, long-chains, shared coordination lead to the formation of complex aggregates (Fig. 6 c, d). The aggregation of molecules as long vs. short networks is analyzed by calculating the number and size of clusters during the production run for last 90 ns of simulation time. A cut off of 2.3 Å, chosen from RDF, is set as the criteria for cluster formation. We define a cluster when one or more atoms of a G4 or TFSI are within 2.3 Å to a Li+ ion. We analyzed these networks as two different ways - individual and shared. For individual networks, mutually exclusive Li…TFSI and Li…G4 clusters are calculated (Table 4). For shared networks, every possible cluster involving Li or G4 or TFSI within 2.3 Å are calculated (Table 5). For these networks, we look at the two extreme conditions : the size of the single largest network and the number of smallest cluster (of size 1, which is essentially an uncoordinated G4 or TFSI). For 1:4 mixture at 300 K, the number of molecules that do not form any cluster is nearly equal to 6 (out of 108) which are mostly TFSI anions. This suggests that in case of excessive solvation with G4, the free uncoordinated molecules are TFSI. For concentrated systems (2:1), 11 (out of 432) TFSI and 4 (out of 216) G4 molecules are left uncoordinated. The analysis of all (individual+shared) clusters shows that 14% TFSI or G4 (total 77.6) are uncoordinated and free at 1:4 at 300 K. As the system is more concentrated, this number first increases for 1:2 to 22% and then decreases for more concentrated systems, 18% for 1:1, and less than 1% for 2:1. The largest networks at various concentrations, which involve the majority of TFSI, G4 forming these via a shared coordination with Li+ ions, are shown in Fig. 7 (a - d).
We obtained some of the prominent finite geometries from these systems (Fig. 7 e - i), which unravel the difference in the degree of correlation in and of TFSI vs. G4. The key arguments are as follows with the index of subfigures:
(e) For 1:4 system, shared clusters mostly involving multiple G4 around a Li+ ion are formed; since only one or rarely two O(G4) coordinate with a Li+ ion, the rest O(G4) coordinate with other Li+ ions and thus forming a long …(G4)..Li…(G4)…Li… network, leaving no individual G4 uncoordinated.
(f) For 1:2 system, some G4 bidentate with a Li+ ion; the other O(G4) atoms form extended networks similar to 1:4.
(g) For 1:1 system, some Li+ ions form close-contact aggregates, which have two Li+ ions bidented by two G4 molecules.
(h) shows single unshared complex of G4 with a Li+ ions; such complexes rarely form shared networks, except at high concentrations ( 1:1), where anions dentate additionally raising the coordination number from 4 to 5 or 6.
(i) shows a Li…TFSI cluster which forms only via shared coordination from multiple TFSI. A single TFSI cannot form a complex like (h) due to rigidity and other geometrical constraints.
It is important to notice here that G4 can form individual unshared complexes with a Li+ ion in a polydentate manner, which TFSI cannot. Thus, a rotation of Li…G4 complex can occur without the need of breaking the coordination, while in case of Li…TFSI, break in coordination is essential for any rotation of an individual TFSI. This explains the poor correlation between and for G4, and highly correlated those for TFSI. A longer residence time only restrict the rotation when there is coordination from multiple similar entities. A polydentate anion or solvent coordinating in an unshared manner would not affect its rotation with its residence time and thus can live longer while rotating faster. The correlation observed here is potentially an important descriptor, that can be further tested in extending the general theories for concentrated electrolytesKrucker-Velasquez and Swan (2021); Dinpajooh et al. (2024) to those with long-chain entities.
IV Conclusion
This work demonstrates the impact of varying salt concentrations and temperature on the structure, dynamics and interplay of residence vs. rotational relaxation times in LiTFSI/G4 mixtures with classical MD simulations. The structural analysis shows that lithium ions preferentially coordinate with glyme molecules, forming [Li(G4)]+ complexes, complementing the earlier experimental and computational studies. Temperature, in the range of 300 K to 500 K, did not significantly affect the local solvation structure of Li+ ion, suggesting that the coordination environment remains stable even at higher temperatures. The dynamic properties indicated that the self-diffusion coefficients of Li+ ions and solvent molecules decreased with increasing salt concentration, resulting in a sluggish movement of ions. This suggests that higher salt concentrations increase the viscosity of the system, thus restricting the overall ionic mobility. At higher salt concentrations, a larger network of clusters is formed, which causes sluggish dynamics of the ions and molecules in the system.
The ability of G4 to form individual unshared polydentate complexes with Li+ allows rotation without breaking coordination, while TFSI must disrupt coordination to rotate. This difference accounts for the varying correlations between rotation and residence time in both cases. A longer residence time restricts rotation only when multiple similar entities coordinate, while a polydentate species with unshared coordination can maintain rapid rotation despite prolonged residence. An experimental examination of this effect could provide valuable insights into the viscosity, ionic conductivity, and thermal stability of LIBs.
Acknowledgment
RR acknowledges funding support from the Department of Science and Technology’s National Supercomputing Mission grant DST/NSM/R&D_HPC_Applications/2021/30. The authors acknowledge the Param Shivay (IIT BHU) and Param Ananta (IIT Gandhinagar) supercomputing facilities for the simulations reported in this work.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.