### v4

# How individual vs shared coordination governs the degree of correlation in rotational vs residence times in a high-viscosity lithium electrolyte

Abstract
Commercially used carbonate-based electrolytes in lithium-ion batteries are susceptible to many challenges, including flammability, volatility, and lower thermal stability compared to LiTFSI-glyme based electrolytes. Mixtures of LiTFSI salt (lithium bis(trifluoromethylsulfonyl)-amide) and glyme-based solvents are potential alternative candidates for commonly used electrolytes. We perform classical molecular dynamics (MD) simulations to study the effect of concentration and temperature on the translational and rotational dynamics. The radial distribution function shows stronger coordination of Li+ ions with tetraglyme(G4), as shown in earlier studies, and forms a stable [Li(G4)]+ cation complex. The self-diffusion coefficients are lower than the values experimentally observed but show improvement over other classical force fields, without charge scaling. An increase in the salt concentrations leads to a higher viscosity of the system and reduces the overall ionic mobility of Li+ ions. Diluting the system with a larger number of G4 molecules leads to shorter rotational relaxation times for both TFSI and G4. Ion-residence times show that Li+ ions form stable and long-lasting complexes with G4 molecules than TFSI anions. The residence time of [Li(G4)]+ complex increases in highly concentrated system due to the availability of fewer G4 molecules to coordinate with a Li+ ion. G4 is also seen to form polydentate complexes with Li+ ions without a shared coordination, allowing rotation without breaking coordination, unlike TFSI, which requires coordination disruption for rotation. This distinction explains the poor correlation between rotation and residence time for G4 and the strong correlation for TFSI.
[…]
Chemistry and Chemical Engineering, California Institute of Technology, Pasadena, California, 91125 United States
1 Introduction
The growing demand for clean energy across the world requires efficient energy storage devices. Batteries and supercapacitors are among the widely used electrochemical devices for energy storage. Lithium-ion batteries (LIBs) have made significant progress in the past three decades in reducing their overall cost with an increase in energy density. The cost has decreased from \$1,000 per kWh in 2009 to \$180 per kWh in 2020, while the energy density has increased from 125 Wh/kg to 250 Wh/kg. The LIB market was \$36.7 billion in 2019 and is expected to reach \$128.3 billion by 2027, with an estimated compound annual growth rate of 18% from 2020 to 2027.feng2022review, nitta2015li. The electrolyte is an important component of batteries that determines the cyclic efficiency and the charge-discharge rate of a battery. Generally, ideal electrolytes should possess high ionic conductivity and good electrochemical and thermal stability. Commercialized LIBs use electrolytes of fluoride salts and organic solvents, which provide high ionic conductivity but are susceptible to leakages and flammability and are responsible for most fire accidents in electric vehicles. Various alternative electrolyte systems, such as ionic liquids (ILs),RANA2023116340 solvate ionic liquids (SILs),Nachaki2024 polymer electrolytes,wu2024fire localized highly concentrated electrolytes,van2023comparative inorganic solid electrolytes,Yu2025 soft-solid electrolytes,Prakash2023 and composite electrolytes,ZHANG2023100316 have been proposed to increase the safety aspects and energy density of the batteries. SILs are considered an alternative to conventional organic electrolytes owing to their unique structural characteristics and higher ionic conductivity. SILs generally formed in equimolar mixtures when a cation (e.g., Li+ here) forms a coordinated complex ([Li(G4)]+) with the donor sites of the solvent molecule, with anions being solvent-separated and occupying the interstitial volume. SILs have higher viscosity compared to organic electrolytes, still they possess higher ionic conductivity which is attributed to the higher concentration of Li(G4)[TFSI].tamura2010physicochemical, di2022glyme The glyme family of solvents, specifically those with higher chain lengths, are a better alternative to organic solvents in ILs due to their lower flammability, less toxicity, and higher thermal stability. In ILs, the currently used higher chain length solvents are triglyme (G3), tetraglyme (G4), and pentaglyme (G5) with LiTFSI and LiPF6 salts. LiTFSI salt shows excellent thermal and chemical stability up to 360 °C and dissociates easily in various types of solvents.wang2019review, DiLecce2022, lingua2024solvate, rajendran2024electrochemical, dong2018charge
Higher chain-length glyme molecules have better physiochemical properties compared to shorter chain-length glyme molecules. Equimolar concentrations of triglyme (G3) and tetraglyme (G4) have been investigated with lithium salts using various experimental (Raman spectroscopy, Small angle X-Ray Scattering, Diffusion nuclear magnetic resonance) and computational techniques.aguilera2015structural, yoshida2011change, tamura2010physicochemical Structural properties like radial distribution function (RDF) and radius of gyration ($R_{\text{g}}$) are useful in predicting the local solvation shell and compactness of atoms or molecules, respectively. The local structure of Li+ ions prefers to remain in a five-fold (or four-fold) coordination with oxygen of G4 and TFSI combined, where the major contribution comes from G4.tsuzuki2014structures, callsen2017solvation, sun2018insight Based on G4 or TFSI coordination, the LiTFSI/G4 complexes are classified as contact ion pairs (CIP) and solvent-separated ion pairs (SSIP). In CIP, oxygen atoms on TFSI directly coordinate with solvated Li+ ions, whereas in SSIP, TFSI anions interact with the [Li(G4)]+ complex. The SSIP configurations were found to be energetically more favorable compared to CIP, stating that Li+ ions prefer to remain as a [Li(G4)]+ complex in the system, also supported by the diffusivity ratio ($D_{\text{Li}^{+}}/D_{\text{G4}}$), which is equal to one for equimolar systems.aguilera2015structural, zhang2014chelate, ho2023understanding
Experimental techniques like 2D IR spectroscopy, along with molecular mechanics-based non-reactive classical molecular dynamics (MD) simulations and first principle calculations, have been used to assess and characterize the distribution of these various structural complexes.rajput2015coupling, callsen2017solvation, rushing2019effect, galvez2021ion, chen2024dominant, tan2024decoding However, the challenge with classical MD is that it does not capture the structural changes in solvated structures during the dynamics very well compared to Ab initio (AIMD) simulations. Still, inaccessibility to longer timescales in AIMD simulations forces the use of classical models, especially for longer time-correlated properties like translational and rotational diffusion. Recently, machine learning interatomic potentials (MLIPs) have emerged as an efficient alternative to both classical and ab initio molecular dynamics for liquid electrolytes simulations in batteries. These models can achieve near–first-principles accuracy while dealing with large system sizes and longer timescales accessible to classical MD.magduau2023machine, goodwin2024transferability, gong2025predictive, levine2025open
Experimentally, the diffusion coefficients for Li+ , TFSI, and G4 are calculated at 30 °C and 80 °C (discussed later in the subsection Translational dynamics).tamura2010physicochemical, yoshida2011change, zhang2014chelate, harte2023accelerated Using MD simulations with a polarizable force field, Dong et al.dong2018charge found that the diffusion coefficients of species at 30 °C and 80 °C are one or two order of magnitude less than the experimental diffusion coefficients. Other research groups also used non-polarizable OPLS-force fields with scaled charges (to include polarization effects) and found similar diffusion coefficients at 227 °C.thum2020solvate No experimental study has been done so far that highlights the system dynamics at such elevated temperaturesharte2023accelerated. The use of scaled charges can model the structure right by over-simplification of the many-body interactions. Thus, the use of more robust methods to compute the partial charge of molecules is required. Dong’s group also calculated the lifetime of the Li+ ions with TFSI and G4 at 373 K, which depends upon the salt concentration. A smaller concentration of salt leads to a greater number of free glyme molecules in the system, which goes through a faster exchange event with glyme molecules binding with Li+ ions, thus reducing the lifetime of the $\text{Li-O}_{\text{G4}}$.dong2018charge The translation diffusion, exchange events, and residence times also dictate the coiling and bending of geometries of G4 and TFSI, respectively. However, no modeling of rotational dynamics has been conducted yet that shows the effect of the coiling/bending of one molecule on another or on the rotational lifetime of molecules. The rotational contribution is an important factor in deciding the lifetime of big continuous clusters of Li+ ions with TFSI and G4. Moreover, the effect of correlated motions on the ionic mobility is not very well understood even in some of the recent studies,mitra2024exploring, philippi2025, bhabani2025 which motivates us to investigate if there is an interplay between residence and rotational lifetimes in such high-viscosity lithium electrolytes. Although, the absolute timescales might be overestimated due to simulation limitations, the observed relative trends and the correlation between residence and rotational times are expected to be reliable and physically meaningful.
Figure 1: (a) Schematic representation of LiTFSI and G4 with partial atomic charges, (b) 1:4, (c) 1:2, (d) 1:1 and (e) 2:1 ratio mixtures of LiTFSI/G4 at 300 K after an equilibration of 10 ns.
This study investigates the extreme range of LiTFSI and G4 mixture compositions (from 1:4 to 2:1) at various temperatures with classical MD simulations using a modified OPLS force-field.jorgensen1996development The partial charges for atoms of LiTFSI and G4 are reparametrized here (Fig. 1). A correct structure prediction and long timescale simulations performed in the present work provide a reliable set of results for diffusion coefficients, residence times, rotational relaxation times, and cluster sizes and theorize a not-yet-explored interplay of these.
2 Computational Details
Four mixtures with different LiTFSI/G4 ratios (1:4, 1:2, 1:1, and 2:1) were studied, as shown in Table S1 of supplementary information. These mixtures were modeled using a charge-reparameterized OPLS force field. Partial atomic charges for the LiTFSI and G4 molecules were calculated (Fig. 1a) using a range-separated hybrid functional wB97x-D4 functionalwb97, GD4 with a sufficiently large basis set def2-tzvpp, utilizing the ORCA 5.0.1 packageORCA5. Partial atomic charges were derived for all atoms of TFSI and G4 from their geometries in Li-bonded states using CHELPG methodchelpg. A fixed charge of +0.87 was assigned to Li+ ions based on the bonded state of the cation and free cation. The PMEessmann1995smooth algorithm was used to compute the long-range electrostatic interactions. The cut-off for Lennard-Jones and electrostatic interactions is set to 10 Å. A time-step of 1 femtosecond is used throughout the simulations. All simulations were conducted using the Gromacs 2021 gromacs software package. For post-processing and analysis, VMD 1.9.3 humphrey1996vmd, Travisbrehm2020travis, and in-house scripts developed in python were used.
The different ratio mixtures were initially minimized using the conjugate-gradient algorithm. Each system was heated from 300 K to 600 K in the NPT ensemble with a heating rate of 50 K per 5 ns to mix it well, followed by an equilibration run in the NVT ensemble for 10 ns at each particular temperature. Fig. 1(b-e) shows post-equilibrated mixtures at 300 K. Velocity rescaling methodbussi2007canonical and Berendsen barostatberendsen1984molecular were used to manage fluctuations in temperature and pressure, respectively. Following this, the systems were cooled again from 600 K to 300 K in the NPT ensemble at the interval of 50 K per 5 ns with 10 ns of NVT equilibration at each particular temperature. Final production runs were performed at 300 K and 500 K in NVT ensemble conditions for 100 ns to obtain sufficiently long trajectories for analysis.
3 Results and discussion
3.1 Solvation structure
(a) (a)
(b) (b)
(c) (c)
(d) (d)
Figure 2: Radial distribution function for (a) Li…O(TFSI), (b) Li…O(G4) at 300 K, and (c) Li…O(TFSI), (d) Li…O(G4) at 500 K for different compositions.
The interatomic interactions of TFSI and G4 around Li+ ions are calculated as radial distribution functions (RDF):
$g(r)=\frac{n(r)}{\rho 4\pi r^{2}\Delta r}$ (1)
$n(r^{{}^{\prime}})=4\pi\rho\int_{0}^{r^{{}^{\prime}}}g(r)r^{2}\,dr\$(2)
where $g(r)$ is RDF, $n(r)$ is the number of particles located at the distance between $r$ and $r+\Delta r$ from the reference particle, $\rho$ is the bulk number density, and $r^{{}^{\prime}}$ is the distance of the first minima of $g(r)$ from each atom. The coordination number $n(r^{{}^{\prime}})$ is calculated by integrating the RDF over the chosen spherical shell around the central atom. The neutron scattering confirms that in an equimolar mixture of LiTFSI and G4, 95 % G4 molecules are coordinated by Li+ ions and the remaining 5 % have an incomplete Li…O(G4) coordination.murphy2016bulk The experimental $g(r)$ peaks are observed at 1.9 Å for Li…O(G4), and 2.0 Å for Li…O(TFSI) in an equimolar (1:1) mixture.murphy2016bulk Our simulations are in a good agreement with these observations as the first solvation shell of a Li+ ion is seen to form at 1.9 Å with O atoms of G4 and at 2.0 Å with O atoms of TFSI (Fig. 2). From the integration of these RDFs (within a cut-off radius of 2.3 Å), the coordination number in the equimolar ratio at 300 K is 3.6 for Li…O(G4) and $\simeq$ 0.7 for Li…O(TFSI). The variation in the coordination number of Li+ ions is observed at different salt concentrations. With the increase in the concentration of glyme molecules, the coordination number of Li…O(G4) increases nearly to 4 at 1:2 and 1:4 ratios. Also, the increase in the concentration of solvent leads to an increase in the intensity of RDF of the Li…O(G4), whereas the intensity of RDF of Li…O(TFSI) decreases as more number of Li+ ions find free G4 and get attached to them. Changing the concentration of salt and solvent does not affect the radius of the first coordination shell around lithium.
The coordination number of Li…O(G4) in equimolar ratio decreased to 3.3 (500 K) from 3.6 (300 K). This could happen due to the accelerated dynamics, which leads to a faster exchange between G4 molecules at higher temperatures within the solvation shell of lithium. This also leads to a decrease in the lifetime of Li…O(G4) complex which is confirmed through ion-pair/ion-solvent existence autocorrelation function (discussed vide infra in subsection Ion-residence times). Also, the intensity of RDF peaks showed a small decrease at 500 K for both Li…O(TFSI) and Li…O(G4) when compared to RDF at 300 K.
3.2 Radius of gyration
Since both the TFSI anion and G4 are long molecular structures, their coiling leads to a significant variation in the radii of gyration ($R_{\text{g}}$). The $R_{\text{g}}$ of these species is an important factor that is sensitive to the temperature and concentration of electrolyte mixture. We used MD trajectories to calculate $R_{\text{g}}$ for TFSI anions and G4 molecules for each ratio and at 300 K and 500 K (Fig. S2 and S3 in supplementary information). As shown in Table 1, the $R_{\text{g}}$ of TFSI and G4 in the most diluted system (1:4) at 300 K is, 0.256 nm and 0.381 nm, respectively. At a higher concentration, 2:1, the $R_{\text{g}}$ for both TFSI and G4 decreases to 0.254 nm and 0.363 nm, suggesting a more coiled configuration, particularly for G4, which shrinks by  5 %. The $R_{\text{g}}$ for both TFSI and G4 decrease in highly concentrated system. This indicates that at higher salt concentrations, most G4 molecules coordinate individually with the Li+ ions, thus increasing their overall compactness. In diluted system, there are more free G4 molecules available to share the coordination leading to higher $R_{\text{g}}$. As expected at higher temperature, $R_{\text{g}}$ for TFSI and G4 increase from 0.256 nm to 0.257 nm and from 0.381 nm to 0.387 nm respectively, when increasing the temperature from 300 K to 500 K. Even these slight changes in $R_{\text{g}}$ is an important factor which can lead to an interplay in translational and rotational dynamics and ion-residence times in these mixtures (discussed vide infra).
Table 1: Average and standard deviation of radius of gyration, $R_{\text{g}}$(nm), for 1:4 and 2:1 system
Temp (300 K) Temp (500 K)
Species Average Std. dev. Average Std. dev.
1:4 TFSI 0.256 0.0006 0.257 0.0006
G4 0.381 0.0016 0.389 0.0018
2:1 TFSI 0.254 0.0002 0.255 0.0003
G4 0.363 0.0007 0.367 0.0034
3.3 Translational dynamics
Figure 3: Trajectory of Li+ ions over a time of 100 ns for different LiTFSI/G4 ratios (a) 2:1 at 300 K, (b) 2:1 at 500 K, (c) 1:4 at 300 K, (d) 1:4 at 500 K.
Table 2: Comparison of self-diffusion coefficients, $D$ (in $10^{-7}$ $\mathrm{cm^{2}/s}$), for equimolar (1:1) LiTFSI/G4 electrolyte reported in various experimental and MD simulations works
Temp (K)